For a process, relation between temperature and volume is
TV3= Constant. I f a monoatomic gas follows this process, find molar specific heat for this process.

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

Assuming the gas to be an ideal gas, this relation is not possible for a monoatomic gas.

This is because ifTV^3=constant,

(PV)V^3=constant
\implies PV^4=constant

Hence\gamma=4

However this value is not true for a monoatomic gas.
Only by knowing that the gas is monoatomic, we may know its molar heat capacity

C_P=(\dfrac{f}{2}+1)R=\dfrac{5}{2}R